\begin{thebibliography}{}

\bibitem[Angeletos and Werning, 2006]{Ang06}
Angeletos, G.-M. and Werning, I. (2006).
\newblock Crises and prices: Information aggregation, multiplicity, and
  volatility.
\newblock {\em {American Economic Review}}, 96(5):1720--1736.

\bibitem[Atkeson, 2000]{Atk01}
Atkeson, A. (2000).
\newblock Discussion on morris and shin.
\newblock In {\em NBER Macroeconomics Annual}, pages 161--170. ed. Bernanke,
  Ben S. and Kenneth Rogoff, Cambridge, MIT Press.

\bibitem[Gorelkina and Kuhle, 2018]{Gor18}
Gorelkina, O. and Kuhle, W. (2018).
\newblock Information aggregation through stock prices and the cost of capital.
\newblock {\em Journal of Institutional and Theoretical Economics},
  174(2):399--420.

\bibitem[Grossman and Stiglitz, 1976]{Gro76}
Grossman, S. and Stiglitz, J. (1976).
\newblock Information and competitive price systems.
\newblock {\em {American Economic Review Papers and Proceedings}},
  66(2):246--253.

\bibitem[Grossman and Stiglitz, 1980]{Gro80}
Grossman, S. and Stiglitz, J. (1980).
\newblock On the impossibility of informationally efficient markets.
\newblock {\em {American Economic Review}}, 70(3):393--408.

\bibitem[Kandel and Pearson, 1995]{Kan95}
Kandel, E. and Pearson, N.~D. (1995).
\newblock Differential interpretation of public signals and trade in
  speculative markets.
\newblock {\em Journal of Political Economy}, 103(4):831--872.

\bibitem[Kuhle, 2016]{Kuh16Priors}
Kuhle, W. (2016).
\newblock A global game with heterogenous priors.
\newblock {\em {Economic Theory Bulletin}}, 4:167--185.

\bibitem[Raiffa and Schlaifer, 2000]{Rai61}
Raiffa, H. and Schlaifer, R. (2000).
\newblock {\em Applied Statistical Decision Theory}.
\newblock Wiley Classics Library.

\bibitem[Vives, 2008]{Viv08}
Vives, X. (2008).
\newblock {\em Information and Learning in Markets}.
\newblock Princeton University Press.

\end{thebibliography}

\documentclass[12pt,a4paper,normalheadings,headsepline,headinclude,bibtotoc,fleqn,review]{article}
\usepackage[latin1]{inputenc}
\usepackage{amsmath}
\usepackage{booktabs}
\usepackage{array}
\usepackage{amsthm}
\usepackage{dcolumn}
\usepackage{endnotes}
\usepackage{units}
\usepackage{marvosym}
\usepackage[lines=10]{geometry}
\usepackage{longtable}
\usepackage{rotating}
\usepackage{expdlist}
\usepackage{geometry}
\usepackage{floatrow}
\usepackage{pgfplots}
\pgfmathdeclarefunction{gauss}{3}{
	\pgfmathparse{1/(#3*sqrt(2*pi))*exp(-((#1-#2)^2)/(2*#3^2))}%
}
\usetikzlibrary{decorations.pathreplacing,calligraphy,backgrounds}
\pgfplotsset{compat=1.16}

\pgfmathsetmacro\valueA{gauss(2.25, 2.25, 0.7)} 
\newtheorem{remark}{Remark}

\usepackage{wrapfig}
\usepackage{expdlist}
\usepackage{amssymb}
\usepackage{gloss}
\usepackage{nomencl}
\usepackage{natbib}

\newtheorem{prop}{Proposition}
\newtheorem{coro}{Corollary}

    \makegloss

\renewcommand{\thesection}{\arabic{section}}

\usepackage{fancyhdr} 
\begin{document}



\begin{center}\huge{Information Aggregation in Markets \\with\\ Analysts, Experts, and Chatbots}
\end{center}

\begin{center} \textit{Wolfgang Kuhle}\\ \textit{Corvinus University of Budapest, Hungary, E-mail wkuhle@gmx.de\\ MEA, Max Planck Institute for Social Law and Social Policy, Munich, Germany}\end{center} 

\noindent\emph{\textbf{Abstract:} The present paper shows that it can be advantageous for traders to publish their information on the true value of an asset even if they (i) cannot build a position in the asset prior to the publication of their information and (ii) cannot charge for the provision of information. The model also shows that the informational content of prices is U-shaped in the number of traders who publish their information. Put differently, information aggregation works best if either no trader, or if every trader publishes his information. Small groups of distinguished experts are, on the contrary, an obstacle to information aggregation. The model's key assumption is that the perception/interpretation of a given piece of published information differs slightly across traders.}\\
\textbf{Keywords: Information Aggregation, Experts, Adverse Information Selection}\\



\noindent\textit{You put two economists in a room, you get two opinions, unless one of them is Lord Keynes, in which case you get three opinions.} Attributed to Winston Churchill
\section{Introduction}\label{Introduction}

It is well known that traders do not share their insights with the public, for if they did, they would loose their edge. In the present paper, we study the conditions under which the opposite is true. That is, we show that the sharing of genuine research information can be advantageous for analysts, even if (i) they did not build a position in the asset prior to the publication of their information and (ii) they cannot charge for the information they publish. 

The model's key assumption is that each traders' understanding of a given published news item, differs slightly. That is, while traders are rational Bayesian agents, each trader who reads an analyst report will interpret it sightly differently.\footnote{Using data on trading volumes, \citet{Kan95} have shown that public announcements, such as earnings reports, are indeed interpreted differently by different traders. To see the same thing, one may consider the heterogeneous interpretation of e.g. interest rate decisions by central banks: The announcement of an interest rate cut by the central bank may be viewed as a positive signal for asset prices by some traders. At the same time other traders may view it as a signal of a weaker economy, which is bad for asset prices.}

We build our model gradually. First, using the model of \citet{Gro76}, we establish a benchmark model of efficient information aggregation in Section \ref{Section Model Perfect Information Aggregation}. This baseline model assumes that the error terms in traders' private signals are uncorrelated, such that equilibrium prices fully reveal the asset's true fundamental.

In Section \ref{Additional Information y Model}, we extend our baseline model, and assume that traders receive an additional private signal over the assets' true value. This second signal, however, contains a common error term, which hinders information aggregation. That is, rational traders have an incentive to use this second piece of information to improve their private understanding of the assets' true value. At the same time, however, the demand functions that traders submit to the market, are now correlated such that equilibrium prices can no-longer aggregate private information efficiently. That is, from the perspective of information aggregation, there is adverse selection in the market regarding the information upon which traders base their asset demands.     

Section \ref{Incentive to Publish Information} builds on Sections \ref{Section Model Perfect Information Aggregation} and \ref{Additional Information y Model}, to show that an individual trader, who publishes his private information on the asset's true value, gains an informational advantage over all other traders, who (i) receive the news item and (ii) who do keep their own private information to themselves. In particular, if only one traders publishes his information, he can fully infer the assets' true fundamental, while the traders, who received the research note can no-longer infer the asset's true value. Finally, we show that the informational content of prices is U-shaped in the number of published research notes.

\textit{Related Literature:} 
Stock prices as aggregators of information are used in many applications.\footnote{See e.g. \citet{Viv08} for a review of this literature.} For example, \citet{Atk01}, \cite{Ang06}, \citet{Kuh16Priors}, \citet{Gor18} and others argue that prices are public signals, which can trigger speculative attacks on banks and currencies. One ``key result" within this literature, as \cite{Ang06}, p. 1732, conclude, is ``that the precision of endogenous public information increases with the precision of exogenous private information. This feature is likely very robust..." Taking this view, the present paper presents a counter example, where the informational content of prices decreases while individual traders' private information improves. 

The rest of the paper is organized as follows. Section \ref{Section Model Perfect Information Aggregation} recalls the \citet{Gro76} model of information aggregation, where the cross-section of agents' private signals reveals the assets' true value. Section \ref{Additional Information y Model} studies the conditions under which the informational content of prices decreases once agents receive two private signals. Section \ref{Incentive to Publish Information} we examine traders' incentive to publish private information. In particular, we show that traders have an incentive to publish their private information, even if they do not get paid for it. Section \ref{Remarks and Extensions} discusses a variation of our model, where, instead of Analysts and Experts, traders turn to chatbots for financial advice. Section \ref{Conclusion} concludes.

\section{A Model of Perfect Information Aggregation}\label{Section Model Perfect Information Aggregation}
Following \citet{Gro76}, we assume that there is an aggregate supply of $K$ shares, each offering an unobservable return:
\begin{eqnarray} \tilde{\theta}=\theta+\eta, \quad \eta\sim\mathcal{N}(0,\sigma_{\eta}). \label{return}\end{eqnarray}
Where the asset's return $\tilde{\theta}$ consists of two independent components $\theta$ and $\eta$. 

There exist a large number of agents, indexed by $i\in[0,1]$. These agents hold a uniform, uninformative, prior over $\theta$. Moreover, each agent receives a private signal regarding the return $\theta$:
\begin{eqnarray}  x_i=\theta+\sigma_x\xi_i, \quad \xi_i\sim\mathcal{N}(0,1). \label{signalx} \end{eqnarray}
where $\xi_i$ is independently distributed across agents, and independent of $\theta$ and $\eta$. Given their information, agents maximize expected utility:
\begin{eqnarray} U_i=-E[e^{-\gamma(k_i(\tilde{\theta}-P))}|x_i,P].\label{Utility} \end{eqnarray}
To solve for the market's equilibrium, we guess and verify that prices are given by a linear function:
\begin{eqnarray} P=b\theta+c. \label{guess1} \end{eqnarray}
Considering (\ref{guess1}), the market price is informationally equivalent to a signal $Z$, which fully reveals the fundamental $\theta$ to traders:
\begin{eqnarray} Z:=\frac{P-c}{b}=\theta.  \label{Zsignalfullyrevealing}\end{eqnarray}
Given utility (\ref{Utility}), traders' demand is:
\begin{eqnarray} k_i=\frac{E[\tilde{\theta}|x_i,P]-P}{\gamma Var(\tilde{\theta}|x_i,P)}=\frac{E[\theta|x_i,Z]-P}{\gamma Var(\tilde{\theta}|x_i,Z)}. \label{demand1} \end{eqnarray}
As mentioned above, the price signal $Z$ is fully revealing $\theta$ such that $E[\tilde{\theta}|x_i,P]=\theta+E[\eta|x_i,P]=\theta$ and $Var(\tilde{\theta}|x_i,P)=\sigma^2_{\eta}$. Demand (\ref{demand1}) thus rewrites as:\footnote{That is, if the Walrasian auctioneer initially chose a random, uninformative, \textbf{non-equilibrium, price}, agents would use their individual signals, and demand would be $k_i=\frac{E[\tilde{\theta}|x_i]-P}{\gamma Var(\tilde{\theta}|x_i)}=\frac{x_i-P}{\gamma (\sigma_x^2+\sigma_{\eta}^2)}$, yielding a market clearing price $P=\theta-\gamma (\sigma_x^2+\sigma_{\eta}^2)$. This price, however, is once again fully revealing, and, in turn, agents would deviate from their demand schedule $k_i=\frac{E[\tilde{\theta}|x_i]-P}{\gamma Var(\tilde{\theta}|x_i)}=\frac{x_i-P}{\gamma (\sigma_x^2+\sigma_{\eta}^2)}$. That is, the use of private information off the equilibrium path, ensures that the Walrasian auctioneer must choose the equilibrium price such that it is fully revealing as in \citet{Gro76}.}
\begin{eqnarray} k_i=\frac{\theta-P}{\gamma \sigma_{\eta}^2}. \label{demand2} \end{eqnarray}
Finally, market clearing requires
\begin{eqnarray} \int_0^1k_idi=K, \end{eqnarray} 
which, together with (\ref{demand1}) yields an equilibrium price:
\begin{eqnarray}
P=\theta-\gamma \sigma_{\eta}^2 K. \label{equilibrium} \end{eqnarray} 
Comparing (\ref{equilibrium}) with (\ref{guess1}) we determine the coefficients $b=1$ and $c=-\gamma \sigma_{\eta}^2 K$. Hence, as in \cite{Gro76}, the market price efficiently aggregates traders' collective information regarding the asset's return $\theta$.

\section{Additional Information Sources}\label{Additional Information y Model}
Let us now assume that in addition to their signal $x_i$, agents have access to an analyst report, or an earnings report. Moreover, we assume that each agents' understanding of this report differs slightly, such that each agent's interpretation of this report is:
\begin{eqnarray}  y_i=\theta+\sigma_{\varepsilon}\varepsilon+\sigma_y\tau_i, \quad \tau_i\sim\mathcal{N}(0,1), \quad \varepsilon\sim\mathcal{N}(0,1), \label{Published Signal}  \end{eqnarray}
where $\sigma_{\varepsilon}\varepsilon$ represents the error in the analyst's report over the true fundamental $\theta$, and $\sigma_y\tau_i$ represents the idiosyncratic noise with which each agent $i$ interprets the report's message. Finally, we assume that idiosyncratic errors $\tau_i$ are independently distributed across agents, and uncorrelated with $\theta$, $\eta$ and $\varepsilon$. 

Using their own information $x_i$, and the financial advice/analyst opinion (\ref{Published Signal}), agents maximize utility:
\begin{eqnarray} U_i=-E[e^{-\gamma(k_i(\tilde{\theta}-P))}|x_i,y_i,P]=-e^{-\gamma k_i(E[\tilde{\theta}|x_i,y_i,P]-P)+k_i^2\frac{\gamma^2}{2}Var[\tilde{\theta}|x_i,y_i,P]}, \label{utility1}\end{eqnarray}
such that individual demand is:
\begin{eqnarray} k_i=\frac{E[\tilde{\theta}|x_i,y_i,P]-P}{\gamma Var(\tilde{\theta}|x_i,y_i,P)}=\frac{E[\theta|x_i,y_i,P]-P}{\gamma Var(\tilde{\theta}|x_i,y_i,P)}. \label{demand} \end{eqnarray}
To work with (\ref{demand}) we have to guess and verify a price function:
\begin{eqnarray}  P=a\theta+b\varepsilon+c. \label{guess2}\end{eqnarray}
It follows from (\ref{guess2}) that the price $P$ carries the same information regarding $\theta$ as a signal:
\begin{eqnarray} Z:=\frac{P-c}{a}=\theta+\frac{b}{a}\varepsilon. \label{Zsignal} \end{eqnarray} 
Combining traders' private information with the price signal (\ref{Zsignal}), yields posteriors:\footnote{See e.g \cite{Rai61}, p. 250.}
\begin{eqnarray} &&E[\theta|x_i,y_i,Z]=x_i+\frac{1}{\alpha}((\frac{b}{a})^2-\frac{b}{a}\sigma_{\varepsilon})(y_i-x_i)\sigma_x^2+\frac{1}{\alpha}(\sigma_y^2+\sigma_{\varepsilon}^2-\frac{b}{a}\sigma_{\varepsilon})(Z-x_i)\sigma_x^2 \label{Posterior mean}\\
&&  Var(\tilde{\theta}|x_i,y_i,Z)=\sigma^2_x+\sigma^2_{\eta}-\frac{\sigma_x^4}{\alpha}(b^2-2b\sigma_{\varepsilon}+\sigma^2_{\varepsilon}+\sigma^2_y) \label{Posterior Variance}\\
&& \alpha=(\sigma_x^2+\sigma_{\varepsilon}^2+\sigma_{y}^2)(\sigma_x^2+(\frac{b}{a})^2)-(\sigma_x^2+\sigma_{\varepsilon}\frac{b}{a})(\sigma_x^2+\sigma_{\varepsilon}\frac{b}{a}), \quad \alpha>0, \label{determinant}\end{eqnarray}
where $\alpha$ is the determinant of the, positive definite, variance-covariance matrix of traders' signals regarding $\theta$. To solve for the model's coefficients $a,b,c$ we compute aggregate demand:
\begin{eqnarray} K^D=\int_0^1k_idi=\int_0^1\frac{E[\theta|x_i,y_i,P]-P}{\gamma Var(\tilde{\theta}|x_i,y_i,P)}di=\frac{\theta+\frac{1}{\alpha}(\sigma_y^2+\sigma_{\varepsilon}^2-\frac{b}{a}\sigma_{\varepsilon})\frac{b}{a}\sigma^2_x\varepsilon-P}{\gamma Var(\tilde{\theta}|x_i,y_i,Z)}.\end{eqnarray} Market clearing thus requires:
\begin{eqnarray} K=\frac{\theta+\frac{1}{\alpha}(\sigma_y^2+\sigma_{\varepsilon}^2-\frac{b}{a}\sigma_{\varepsilon})\frac{b}{a}\sigma^2_x\varepsilon-P}{\gamma Var(\tilde{\theta}|x_i,y_i,Z)},  \label{Market clearing} \end{eqnarray}
which solves for the price $P$ such that:
\begin{eqnarray} P=\theta+\frac{1}{\alpha}(\sigma_y^2+\sigma_{\varepsilon}^2-\frac{b}{a}\sigma_{\varepsilon})\frac{b}{a}\sigma^2_x\varepsilon-K\gamma Var(\tilde{\theta}|x_i,y_i,Z).\label{Price2}\end{eqnarray}
Combining (\ref{Price2}) and (\ref{guess2}), while taking into account (\ref{Posterior mean})-(\ref{determinant}), yields the model coefficients $a,b,c$: 
\begin{prop} The model features two equilibria with coefficients $a=1,b=0$, and $a=1,b=\frac{\sigma_x^2\sigma_{\varepsilon}}{\sigma^2_x+\sigma^2_y}$, respectively. \label{Two equilibria Proposition}\end{prop}
\begin{proof}
    First, comparison of (\ref{Price2}) and (\ref{guess2}) idicates that we have $a=1$. Second, the coefficient $b$, is the solution to the equation $b=\frac{1}{\alpha}(\sigma_y^2+\sigma_{\varepsilon}^2-\frac{b}{a}\sigma_{\varepsilon})\frac{b}{a}\sigma^2_x$. Taking into account that $\alpha$ is a function of $b$, we have two solutions: $b=0$ and $b=\frac{\sigma_x^2\sigma_{\varepsilon}}{\sigma^2_x+\sigma^2_y}$. Finally, substituting $a$ and $b$ into (\ref{Posterior Variance}) and (\ref{determinant}), yields $Var(\tilde{\theta}|x_i,y_i,Z)$, and thus the constant $c=-K\gamma Var(\tilde{\theta}|x_i,y_i,Z)$.\end{proof}
\noindent Regarding Proposition \ref{Two equilibria Proposition} we remark:
\begin{remark}
    The equilibrium $b=0$, implies that agents ignore the expert opinion/the analyst report, and the resulting price $P=\theta+c$ efficiently aggregates agents' dispersed private information over the fundamental $\theta$, as in our model of perfect information aggregation in Section \ref{Section Model Perfect Information Aggregation}. This equilibrium is, however, unstable: once a mass $\lambda>0$ of agents rely on the expert's opinion, prices are no-longer fully revealing, and the remaining mass of agents $1-\lambda$ have an incentive to also rely on the expert's opinion. 
\end{remark}
In what follows, we focus on equilibria where agents put a positive weight on the published news item:
\begin{eqnarray} a=1, \quad b=\frac{\sigma_x^2\sigma_{\varepsilon}}{\sigma^2_x+\sigma^2_y}, \quad c=-\gamma K Var(\tilde{\theta}|x_i,y_i,Z). \end{eqnarray}
Such that:
\begin{eqnarray}
    P=\theta+\frac{\sigma_x^2\sigma_{\varepsilon}}{\sigma^2_x+\sigma^2_y}\varepsilon-\gamma K Var(\tilde{\theta}|x_i,y_i,Z). \label{Price3}
\end{eqnarray}
Using (\ref{Price3}), we thus have:
\begin{prop}\label{Prop 0}
    The price reveals the true fundamental $\theta$ with precision $\alpha_z=\frac{1}{b^2}=\alpha_{\varepsilon}(1+\frac{\alpha_x}{\alpha_y})^2$, where $\alpha_x=\frac{1}{\sigma_x^2}$, $\alpha_{\varepsilon}=\frac{1}{\sigma^2_{\varepsilon}}$, and $\alpha_y=\frac{1}{\sigma^2_y}$.
\end{prop}
According to Proposition \ref{Prop 0}, the informational content of prices is increasing in the private signal's precision $\alpha_x$, as well as in the precision of the underlying analyst report $\alpha_{\varepsilon}$. On the contrary, the precision $\alpha_{y}$, with which agents understand the analyst's report, is reducing the quality of the price signal. That is, agents place a greater weight on the analyst report once the precision $\alpha_y$, with which they understand this report, increases. In turn, the weight that the common noise term $\varepsilon$ has in the price function increases, which dilutes the informational content of prices.

Comparing the current model to our baseline model in Section \ref{Section Model Perfect Information Aggregation} yields:
\begin{prop}\label{Prop 01}
    The informational content of prices falls if traders have access to an analyst report. 
\end{prop}
\begin{proof}
    Without analyst reports, the market aggregates information as in Section \ref{Section Model Perfect Information Aggregation}. That is, the equilibrium price in equation (\ref{Zsignalfullyrevealing}) fully reveals $\theta$. On the contrary, once traders read the analyst's report, conveying information as in (\ref{Published Signal}), the new equilibrium price (\ref{Price3}) no-longer reveals $\theta$ fully. 
\end{proof}

\section{Incentive to Publish Information}\label{Incentive to Publish Information}
Suppose now that one trader, indexed by $j$, can choose to publish his information $x_j$ in a news-outlet, or on an investment forum, or he can appear as an expert on television to share his knowledge $x_j$ with the public. 

To align notation with Section \ref{Additional Information y Model}, we relabel the noise on trader $j's$ published signal such that $x_j=\theta+\sigma_x\xi_j$, as $\xi_j=:\varepsilon$ and $\sigma_x=:\sigma_{\varepsilon}$. In turn, the population of traders $i$, who read the published report, receive information: 
\begin{eqnarray}
    y_i=x_j+\sigma_y\tau_i, \quad x_j=\theta+\sigma_{\varepsilon}\varepsilon. \label{Published informarion}
\end{eqnarray}
Once we note that the agent who publishes this information is himself small, such that he does not move prices, other than through the publication of his information, we can compare the models of Sections \ref{Section Model Perfect Information Aggregation} and \ref{Additional Information y Model} to analyze the impact that the publication of a private signal has on the market and its capacity to aggregate information:  
\begin{coro}\label{Prop 1}
    The informational content of prices falls once one trader publishes his information. 
\end{coro}
\begin{proof}
    Follows directly from Proposition \ref{Prop 01}.
\end{proof}
Regarding the incentive to publish information, we note:

\begin{prop}\label{Prop 2}
    The trader $j$, who published his private information, gains an information advantage over all other traders. 
\end{prop} 
\begin{proof}\label{Prop 2}
    If no trader publishes his information, the price signal (\ref{Zsignalfullyrevealing}) fully reveals $\theta$. On the contrary, if one trader $j$ publishes his information (\ref{Published informarion}), the particular trader $j$ can combine his knowledge of $x_j=\theta+\sigma_{\varepsilon}\varepsilon$ with equation (\ref{Price3}), and solve these two linear equations for $\theta$ and $\varepsilon$. The other traders $i$, on the contrary, have to rely on a noisy posterior distribution, with mean (\ref{Posterior mean}) and variance (\ref{Posterior Variance}) to infer $\theta$. Finally, we note that trader $j$ is small (has mass zero) relative to the large market, and thus his (finite) demand $k_j=\frac{E[\theta|x_j,P]-P}{\gamma Var(\tilde{\theta}|x_j,P)}=\frac{\theta-P}{\gamma Var(\eta)}$ does not impact overall equilibrium.
\end{proof}
Finally, if more than one trader publishes his information, we have: 
\begin{prop}\label{Prop 3}
    The informational content of prices is U-shaped in the number of published private research reports. 
\end{prop}
\begin{proof}  
In Proposition \ref{Prop 1}, we showed that the publication of one signal reduces the informational content of prices. In order to prove Proposition \ref{Prop 3}, it thus remains to show that if \textit{all traders} were to publish their information, then the market price would become fully revealing again. Indeed, if all traders publish their private signals in the form of (\ref{Published informarion}), investors can learn $\theta$ just by aggregating the published reports. That is, every investor receives a continuum of reports:
\begin{eqnarray}
    y_i=\theta+\sigma_x\xi_i+\sigma_y\tau_i,\quad i\in[0,1],
\end{eqnarray}
which allow to directly average over these signals, such that, by the strong law of large numbers, we have:
\begin{eqnarray}
    \int_0^1y_idi=\theta.
\end{eqnarray}
Hence traders demands and equilibrium prices are fully revealing, as in Section \ref{Section Model Perfect Information Aggregation} where $k_i=\frac{E[\theta|\mathcal{I}]-P}{\gamma\sigma^2_{\eta}}=\frac{\theta-P}{\gamma\sigma^2_{\eta}},$ such that $P=\theta-\gamma\sigma^2_{\eta}K.$
\end{proof}

\section{Chatbots as Experts}\label{Remarks and Extensions}
Next to analysts and experts, traders may turn to chatbots for financial advice:  
\begin{remark}
    Suppose agents $i\in[0,1]$, submit queries $q_i=f(x_i)$ to an AI chatbot, where $x_i=\theta+\sigma_x\xi_i,$ $\xi_i\sim\mathcal{N}(0,1)$ is the agents' private information, which they use to formulate their questions. Suppose also that these questions can be inverted s.t. $x_i=f^{-1}(q_i)$, then the chat-bot can infer the asset's fundamental $\theta$.\footnote{Indeed, as we have seen in the proof of Proposition \ref{Prop 3}, to learn $\theta$, it suffices that the chatbot infers a noisy signal $ y_i=x_i+\sigma_y\tau_i, i\in[0,1]$ from agents' queries.} Moreover, if the chat-bot rewards traders' queries with answers $AI_i=\theta+\sigma_{\varepsilon}\varepsilon+\sigma_{\tau}\tau_i$, $\varepsilon\sim\mathcal{N}(0,1),\tau_i\sim\mathcal{N}(0,1)$, traders will have an incentive to keep submitting their questions even though these questions reveal their private information $x_i$, which allows the AI chat-bot, and its owners, to free-ride on traders' information. 
\end{remark} 


\section{Conclusion}\label{Conclusion}

\citet{Gro80} argue that prices cannot convey information efficiently, for if they did, traders would not be compensated for the cost of collecting information. The present model, turns this argument on its head: Traders who give their information away gain an informational advantage over their peers who keep their information to themselves. 

There exists a large literature, which emphasizes that markets aggregate dispersed private information.\footnote{See \citet{Atk01}, \cite{Ang06}, \citet{Kuh16Priors}, \citet{Viv08}, or \citet{Gor18} for such applications.} One ``key result," in this literature is, as \cite{Ang06}, p. 1732, stress, ``that the precision of endogenous public information increases with the precision of exogenous private information." The present model provides a counter example to this view, where improvements agents' private information reduce the precision with which prices reveal the asset's fundamental value. That is, the present model provides an example where improvements in agents' private information coincide with a reduction in endogenous public information.

Finally, we found that the informational content of prices is U-shaped in the number of published expert opinions. That is, a small community of distinguished experts is more harmful to information aggregation than a large group of experts, where each individual ``expert" is taken less seriously. Finally, in the ``expert-free limit," where each trader publishes his opinion on the asset's value, prices once again reflect fundamentals accurately.

\addcontentsline{toc}{section}{References}
\markboth{References}{References}
\bibliographystyle{apalike}
\bibliography{References}

\end{document}